OpenStudy (anonymous):

which ordered pair is a solution of the equation y=x-3 A.(-2,5) B.(-5,2) C.(2,5) D.(5,2) its on my home work and im a little confused on it can anyone help me

OpenStudy (anonymous):

plz

pooja195 (pooja195):

plug them into the equation

pooja195 (pooja195):

\[(x,y)\]

OpenStudy (16shuston):

yeah so plug in a set of ordered pairs for example we will take A. (-2,5) so 5=-2-5 and then it will = 5=-7 which is not correct.. so find one that is 5=5

OpenStudy (anonymous):

how did u do that

OpenStudy (16shuston):

x is -2 and 5 is y

OpenStudy (16shuston):

did you get your answer??

OpenStudy (anonymous):

no

OpenStudy (16shuston):

okay lets do b then (-5,2) 2=-5-3 simplify that to 2=-8, they do not equal.. do yo see that?

OpenStudy (16shuston):

*you

OpenStudy (anonymous):

kinda yea

OpenStudy (16shuston):

okay so then do it with c and D and tell me which one you think it is

OpenStudy (anonymous):

got it thx

OpenStudy (16shuston):

no problem you should fan me!!! please :)